OpenStudy (anonymous):

How would I find the radius of a circular cylinder of hieght 10m with the same volume a a 2m by 2m by 10m box?

OpenStudy (anonymous):

Volume box = volume cylinder l*w*h=pi*r^2*h 2*2*10=pi*r^2*10 Then solve for r

OpenStudy (king):

\[r=\sqrt{7}\]

OpenStudy (anonymous):

This is what I got